OpenStudy (anonymous):

Ryan, Jessica, and Mark each solved this system of equations. All of their answers are different. Find and explain the mistakes. What is the correct answer? y = 3x + 2 6x - 2y = 10

OpenStudy (mathstudent55):

y = 3x + 2 6x - 2y = 10 Since the first equation is already solved for y, you can easily use the substitution method. y = 3x + 2, ,so where you see y in the second equation, substitute 3x + 2. 6x - 2y = 10 6x - 2(3x + 2) = 10 6x - 6x - 4 = 10 -4 = 10 Since we get a statement, -4 = 10, that is not true, that means there is no solution for this system of equations.